# The WINTER Observatory: A One-Degree InGaAs Survey Camera to study the Transient Infrared Sky

### v1

@@ III.2 Readout
Custom electronics, firmware, and software are required to control the InGaAs sensors and to accommodate WINTER’s size, heat, and readout speed requirements. The sensors are individually managed by dedicated electronics controlled by FPGAs, sending data over fiber to a computer located in a nearby shed to minimize heat buildup in the dome. This configuration allows for low-noise and low-latency data acquisition within a relatively compact, freely rotating instrument. The custom readout scheme also leverages the fast readout of the sensors. Reading each pixel continuously throughout an exposure–sampling multiple times per second–enables noise reduction through common methods such as Fowler pairing and sample-up-the-ramp. Furthermore, the fast readout software works in tandem with robotic observing to enable near-real-time transient detection and alerts. Further discussion on the WINTER readout design can be found in Malonis:2020 and Frostig:2022SPIE.
@@ III.2.2 Firmware and software
The FPGA firmware runs the sensors, TEC, and focal plane readout electronics locally and interfaces with the WINTER control software. The firmware is written in Verilog, fully modeled in Vivado Simulator, and provides three main functions: electronic control, sensor control, and interfacing with the software. To operate the sensors, the firmware provides clocking, imaging signals, and writes the highly programmable ROIC settings, enabling imaging in various modes at a range of frame rates. Moreover, the firmware controls the power sequencing, telemetry, and housekeeping, and data collection of the sensors and readout electronics along with a safety watchdog overseeing system temperatures and connectivity to the software. Finally, the firmware communicates with the software through a USB 3.0 connection using the Opal Kelly Host Interface, providing an interface to enact commands from the software and to collect and stream sensor data.
The six FPGA modules stream data over USB 3.0, converted to OM4 multimode fiber, to a readout computer in a neighboring shed. Each sensor has its own Icron USB 3.0 Spectra 3022, allowing full bandwidth data link while extending the range of the USB allowing the readout computers to be located outside of the telescope dome and providing electrical isolation of the USB link to avoid ground loops in the readout electronics. An independent multi-threaded, python-based software daemon independently controls each sensor in parallel for rapid data readout and temperature monitoring, and a top level daemon coordinates the image acquisition on each sensor and collates the data into a single multi-extension FITS file.
@@ III.2.3 Read-out modes
The WINTER sensors feature two implemented read-out modes: integrate while read (IWR) and non-destructive read (NDR). IWR is a simpler read out mode where a single image is sampled at the end of the exposure. In this case, the data are read out while the subsequent image is exposing to reduce overheads. This simpler mode was deployed during commissioning of the WINTER sensors and is the current operating mode, since exposures are not read-noise limited.
Alternatively, NDR allows for a reduction in read noise (see Section III.4.4) and is an available mode for observing with WINTER, although it is not currently used in the survey. Below we describe the read-out scheme for NDR.
For short exposures, data are streamed in at seven frames per second, utilizing every available frame, and for longer exposures, the FPGA streams in every other frame, pausing between frames to control temperature deviations.
At the end of an exposure, the continuously sampled frames are immediately pre-processed into single-frame, six-sensor FITS format images which are then downlinked from the observatory for further data reduction (see Section VI.3). To pre-process the WINTER data, the many readout frames are collapsed into one image with one of two techniques: (1) correlated-double sampling (CDS), in which the final frame is subtracted from the first frame, or (2) sample-up-the-ramp (SUTR), in which a linear fit is used to get a slope of the signal in each pixel. SUTR reduces read noise (see Section III.4.4) and allows for recovering data affected by saturation, cosmic rays, and streaming errors, but also adds significant data volumes and overhead time.
To facilitate transient discovery, WINTER requires rapid data pre-processing and reduction to minimize data storage volume and enable real-time decisions on candidate events and observing strategies. The use of graphics processing units (GPUs) is well-suited to efficiently process high-definition imaging data. WINTER’s readout computer has two commercial NVIDIA GeForce RTX 3080 Ti GPUs, each one processing three sensors’ images in parallel. For ease of use, the GPU code uses cupy, an open-source GPU-implemented python library modeled after numpy. Implementing the nonlinearity correction and NDR linear fit on GPUs can lead to a factor of 10 speed up from a CPU-only implementation for a five minute exposure with six sensors.
@@ III.3 Thermal control
WINTER’s InGaAs sensors use a vacuum-sealed package with a two-stage thermoelectric cooler (TEC) at the core of their cooling system. Each sensor has its own mechanical and thermal module which includes a kinematic mounting system for holding the detector allowing manual focus adjustments during assembly, and provides the required cooling to the detectors TEC assembly. An oxygen free high conductivity (OFHC) copper sled, sandwiched between the back of the sensor and the readout electronics, draws heat away from the sensor with two sintered copper/water heat pipes leading to a liquid heat exchanger, placing any liquid outside of the main instrument box and away from the optics and electronics (Frostig:2022SPIE). The heatpipes soldered with solder paste into a OFHC copper flange which is bolted to a small liquid heat exchanger with an indium foil interface material to maximize conductive coupling. An OptiTemp OTC 1.0A closed-loop 70%/30% distilled water/propylene glycol mix chiller located in a shed next to the telescope dome provides parallel cooling to each of the six focal plane modules through a manifold located on the telescope rotator. The chiller is set to 10∘C which maintains the warm side of the TEC to 12∘C through the 15 m of hose between the chiller and the instrument. The sensor readout computer controls the TEC setpoint with a proportional-integral-derivative (PID) loop. The instrument itself is sealed with a series of gaskets to keep out dust, pollen, and other environmental contaminants common at Palomar Observatory. To avoid condensation within the camera enclosure during humid days, a dry air purge system consisting of a 1 horsepower Powerex oil-less compressor and a nano purification systems dryer delivers air with a -40∘C dewpoint. This system delivers 28 liters per minute of air which is piped through the camera with a series of 1/4 inch hoses with small holes drilled in them, which provides sufficient airflow through the box to keep critical electronics like the FPGA modules from overheating.
WINTER is designed for sky-background-limited photometry, with TEC-based cooling to reducing the dark current levels below the sky background. The smaller prototype instrument employed a single-stage TEC and demonstrated background-limited performance between -C to -C, with dark current halving for every 7∘C of cooling (Simcoe:2019aps). However, for the manufactured WINTER sensors, a combination of TEC under-scope and insufficient process control in sensor packaging resulted in decreased cooling capacity and varied final operating temperatures across the six sensors (between -10∘C and -30∘C). Section III.4.3 provides a detailed discussion of the sensor cooling.
@@ III.4 Performance
The WINTER project initiated the first manufacturing run of these sensors, incorporating a new ROIC design. The vendor manufactured the sensors on a best-effort basis and enough sensors were produced to fill in WINTER’s focal plane; however, the sensors did not meet several key performance specifications, as discussed in this section. Overall, this deviation from expected performance results in a factor of 10 reduction in instrument efficiency (2 magnitudes decrease in depths) from the instrument design. In total, fourteen sensors were received, with varying levels of functionality. Six of the highest-performing sensors were integrated into WINTER, with five non-responsive units allocated for thermal testing and three kept as spares.
@@ III.4.1 Conversion gain
Characterizing the WINTER sensors begins with measuring the conversion gain that links photoelectron signal to the corresponding digital output. Based on the ROIC design, the gain is estimated to be g=1.6 e-/DN (see Frostig:thesis for the full calculation). Typically, this value is experimentally verified using a Photon Transfer Curve (PTC) measurement, which assumes the presence of a shot-noise limited regime where noise increases with signal level N as . Multiple PTC approaches explored in Frostig:thesis, reveal nonlinear responses (see Section III.4.5) and differing gain estimates for the NDR and IWR read out modes. Each method produces a gain estimate within a error of the expected g=1.6 e-/DN. Furthermore, the gains vary by sensor, between the eight read out channels within each sensor, and across the pixels within those read out channels.
For simplicity, we present one measurement of the gain here and use it throughout the paper, but note the significant spread in gain by pixel, requiring careful flat-fielding. Figure 4 shows the range of slopes from a PTC for one readout channel of a WINTER sensor, taken with a set of laboratory flats on an integrating sphere. The variance is taken for a stack of ten dark-subtracted frames at each exposure time, with eight rounds of sigma clipping to reject pixels outside of the three standard deviations of the mean. A line is fit to the slope of the variance vs mean signal for each pixel. This analysis predicts a mean gain of g=2.5 e-/DN but shows a spread of responses across pixels within this read out, with some pixels consistent with the designed gain of g=1.6 e-/DN. Given this uncertainty, we present all analyses throughout this study in both measured counts (DN) and converted to electrons (e-).
@@ III.4.2 Quantum efficiency
We measured a quantum efficiency (QE) of less than 10% for the WINTER sensors, significantly lower than the designed 80%, indicating substantial losses in photon-to-electron conversion efficiency. This measurement was validated through three independent tests: (1) laboratory characterization of the full system using a controlled light source, (2) direct laboratory measurement of the sensors’ photocurrent response, and (3) on-sky calibration with reference stars. Laboratory measurements were made using a 20 inch integrating sphere illuminated with a laser driven light source from Energetiq and a series of optical and near infrared narrow bandpass filters. The absolute optical flux was determined using a silicon photodiode with a NIST-traceable responsivity calibration, allowing the quantum efficiency to be measured between 500 and 1050 nm. The results revealed nonlinear behavior (see Section III.4.5) and a peak response of approximately 5% QE, suggesting a manufacturing issue leading to deviations from the expected efficiency (Figure 6) Furthermore, in standard QE and photon transfer curve (PTC) measurements, conversion gain and QE are inherently degenerate, necessitating an independent test to pinpoint the root cause of the discrepancy.
To isolate the sensor photodiode array’s intrinsic response to light from any effects introduced by the readout integrated circuit (ROIC) or the custom readout chain, we conducted a secondary set of QE tests. These tests employed an external source measurement unit to directly supply the sensor bias voltage, allowing precise voltage control and direct photocurrent measurement on the sensor while bypassing the readout circuitry via programmable logic on the ROIC. This approach enabled a direct measurement of photocurrent response, confirming the measurement of a peak QE of (Figure 6). We define the quantum efficiency as the ratio of the number of electrons generated per incident photon. Given an input optical power (in nW) and a measured photocurrent (in µA), the QE can be expressed as:
@@
Finally, on-sky flux observations of standard stars independently measure WINTER’s full instrument and telescope throughput. Accounting for as-measured reflectivities of the mirrors and lenses, the results confirm an average sensor QE of 10 (see Frostig:2024 for a detailed calculation). This analysis includes sensor serial number SCI022, which was previously in the Starboard B position in the full instrument, which has a laboratory measured QE of 9.6% (Figure 6).
The decreased QE when directly probed on the sensors most likely arises from an issue in the manufacturing or quality control during hybridization, photodiode array production, or processing. The test procedures confirms that poor QE does not result from deficiencies in the ROIC or amplification/read-out chain. Since the InGaAs growth and photodiode array design made use of mature processes at Teledyne-FLIR, we conclude that this production lot of a new device experienced issues in preparation and bonding of InGaAs to the Si multiplexer, a hypothesis that could be verified with destructive testing and microscopy of a sensor from the lot.
@@ III.4.3 Dark current
Figure 7 displays the dark current in the six WINTER sensors by temperature in the final integrated camera on the telescope. The sensors reach a minimum temperature between T C and T C, unable to achieve the designed cooling of T C. The temperature discrepancy arises from two key factors. First, the TEC selection process did not include realistic modeling of the sensor packaging, resulting in the use of an undersized two-stage TEC (Laird MS2-107-10-10-12-12-11-W8). Second, the final design placed an excessive heat load on the TEC through parasitic radiative coupling and thermal conduction through wire bonds from the pins. Temperature variations between sensors are influenced by differences in TEC bonding, vacuum seal integrity, and the temperature of each sensor board, which impacts the TEC’s hot-side temperature. The dark current decreases approximately linearly with the logarithm of the temperature down to about C, beyond which it plateaus, possibly due to thermal infrared emission from the detector housing.
@@ III.4.4 Read noise
One key advantage of the WINTER sensors over other InGaAs detectors is the ability to read out the sensors non-destructively throughout an exposure. This allows for read-noise reduction through algorithms like sample-up-the-ramp (NDR), recovery of saturated data, and real-time elimination of cosmic rays (Robberto_2014). The non-destructive read mode is not currently enabled in nightly WINTER operations. This section explores the read noise measured with and without non-destructive read.
To evaluate the read noise in the WINTER sensors we analyze the reference rows of each sensor. These reference rows, positioned outside the active 1920 × 1080 photosensitive region, are insensitive to incident radiation. As a result, any signal variations in these pixels arise predominantly from read noise without incorporating the effects of pixel response nonlinearity (see Section III.4.5) or photon shot noise. Figure 8 illustrates an analysis of the signal standard deviation for each reference pixel over 384 frames in a non-destructive read ramp. The median read noise for a single readout, equivalent to the current on-sky IWR mode, is approximately 24.5 DN RMS (equivalent to 61.3 e-, assuming a gain of 2.5 e-/DN), which is close to the initial project requirement of 45 e- RMS, depending on the exact gain per pixel.
By continuously reading out the sensors with non-destructive sampling, multiple exposures of the same pixel can be averaged before a reset, reducing the random electronic noise in accordance with the theoretical scaling. Figure 8 demonstrates how this technique lowers the effective read noise as frames are binned in increasing groups from two to 64. The standard deviation of the readout signal along a best-fit line decreases, with the lowest read noise reaching approximately 3 DN (7.5 e-) when averaging in groups of 164 frames.

### v2

@@ III.2 Readout
Custom electronics, firmware, and software are required to control the InGaAs sensors and to accommodate WINTER’s size, heat, and readout speed requirements. The sensors are individually managed by dedicated electronics controlled by FPGAs, sending data over fiber to a computer located in a nearby shed to minimize heat buildup in the dome. This configuration allows for low-noise and low-latency data acquisition within a relatively compact, freely rotating instrument. The custom readout scheme also leverages the fast readout of the sensor and works in tandem with robotic observing to enable near-real-time transient detection and alerts. Further discussion on the WINTER readout design can be found in Malonis et al. (2020) and Frostig et al. (2022b).
@@ III.2.2 Firmware and software
The FPGA firmware runs the sensors, TEC, and focal plane readout electronics locally and interfaces with the WINTER control software. The firmware is written in Verilog, fully modeled in Vivado Simulator, and provides three main functions: electronic control, sensor control, and interfacing with the software. To operate the sensors, the firmware provides clocking, imaging signals, and writes the highly programmable ROIC settings, enabling imaging in various modes at a range of frame rates. In addition, the firmware controls power sequencing, telemetry, housekeeping, and data collection of the sensors and readout electronics, along with a safety watchdog that oversees system temperatures and connectivity to the software. Finally, the firmware communicates with the software through a USB 3.0 connection using the Opal Kelly Host Interface, providing an interface to enact commands from the software and to collect and stream sensor data.
The six FPGA modules stream data over USB 3.0, converted to OM4 multimode fiber, to a readout computer in a neighboring shed. Each sensor connects to its own Icron USB 3.0 Spectra 3022 extender, which provides full-bandwidth data transfer, extends the USB range to a remote readout computer outside the telescope dome, and electrically isolates the link to prevent ground loops. An independent multi-threaded, python-based software daemon independently controls each sensor in parallel for rapid data readout and temperature monitoring. A top level daemon coordinates the image acquisition on each sensor and collates the data into a single multi-extension FITS file.
@@ III.2.3 Read-out modes
The WINTER sensors feature two implemented read-out modes: integrate while read (IWR) and non-destructive read (NDR). IWR is a simpler read out mode where a single image is sampled at the end of the exposure. In this case, the data are read out while the subsequent image is exposing to reduce overheads. This simpler mode was deployed during commissioning of the WINTER sensors and is the current operating mode, since exposures are not read-noise limited. Alternatively, NDR allows for a reduction in read noise (see Section III.4.4) and is an available mode for observing with WINTER, although it is not currently used in the survey. Below we describe the read-out scheme for NDR.
For short exposures, data are streamed in at seven frames per second, utilizing every available frame, and for longer exposures, the FPGA streams in every other frame, pausing between frames to control temperature deviations. At the end of an exposure, the continuously sampled frames are immediately pre-processed into single-frame, six-sensor FITS format images which are then downlinked from the observatory for further data reduction (see Section VI.3). To pre-process the WINTER data, the many readout frames are collapsed into one image with one of two techniques: (1) correlated-double sampling (CDS), in which the final frame is subtracted from the first frame, or (2) sample-up-the-ramp (SUTR), in which a linear fit is used to get a slope of the signal in each pixel (Chapman et al., 1990). SUTR reduces read noise and recovers data affected by saturation, cosmic rays, and streaming errors (Robberto, 2014), but also adds significant data volumes and overhead time.
To facilitate transient discovery, WINTER requires rapid data pre-processing and reduction to minimize data storage volume and enable real-time decisions on candidate events and observing strategies. The use of graphics processing units (GPUs) is well-suited to efficiently process high-definition imaging data. WINTER’s readout computer has two commercial NVIDIA GeForce RTX 3080 Ti GPUs, each one processing three sensors’ images in parallel. For ease of use, the GPU code uses cupy, an open-source GPU-implemented python library modeled after numpy. Implementing the nonlinearity correction and NDR linear fit on GPUs leads to a factor of 10 speed up from a CPU-only implementation for a five minute exposure with six sensors.
@@ III.3 Thermal control
WINTER’s InGaAs sensors use a vacuum-sealed package with a two-stage thermoelectric cooler (TEC) at the core of their cooling system. Each sensor has its own mechanical and thermal module which includes a kinematic mounting system for holding the detector. The mount allows for manual focus adjustments during assembly and provides the required cooling to the detectors’ TEC modules. An oxygen free high conductivity (OFHC) copper sled, sandwiched between the back of the sensor and the readout electronics, draws heat away from the sensor with two sintered copper/water heat pipes leading to a liquid heat exchanger, placing any liquid outside of the main instrument box and away from the optics and electronics (Frostig et al., 2022b). The heat pipes are soldered with solder paste into a OFHC copper flange which is bolted to a small liquid heat exchanger with an indium foil interface material to maximize conductive coupling. An OptiTemp OTC 1.0A closed-loop 70%/30% distilled water/propylene glycol mix chiller located in a shed next to the telescope dome provides parallel cooling to each of the six focal plane modules through a manifold located on the telescope rotator. The chiller is set to 10∘C which maintains the warm side of the TEC to 12∘C through the 15 m of hose between the chiller and the instrument. The sensor readout computer controls the TEC setpoint with a proportional-integral-derivative (PID) loop. The instrument itself is sealed with a series of gaskets to keep out dust, pollen, and other environmental contaminants common at Palomar Observatory. To avoid condensation within the camera enclosure during humid days, a dry air purge system consisting of a 1 horsepower Powerex oil-less compressor and a nano purification systems dryer delivers air with a -40∘C dewpoint. This system delivers 28 liters per minute of air which is piped through the camera with a series of 1/4 inch hoses with small holes drilled in them, which provides sufficient airflow through the box to keep critical electronics like the FPGA modules from overheating.
@@ III.4 Performance
The WINTER project commissioned the first batch of these InGaAs sensors, which included designing a new, larger ROIC. The vendor manufactured the sensors on a best-effort basis and enough sensors were produced to fill in WINTER’s focal plane; however, the sensors did not meet several key performance specifications, as discussed in this section. In total, fourteen sensors were received, with varying levels of functionality. Six of the highest-performing sensors were integrated into WINTER, with five non-responsive units allocated for thermal testing and three kept as spares. Overall, this deviation from expected performance results in a factor of 10 reduction in instrument efficiency (2 magnitudes decrease in depths) from the instrument design.
@@ III.4.1 Conversion gain
Characterizing the WINTER sensors begins with measuring the conversion gain that links photoelectron signal (e-) to the corresponding digital output (in digital numbers, DN). Based on the ROIC design, the gain is estimated to be g=1.6 e-/DN (see Frostig (2024) for the full calculation). Typically, this value is experimentally verified using a Photon Transfer Curve (PTC) measurement, which assumes the presence of a shot-noise limited regime where noise increases with signal level N as . Multiple PTC approaches explored in Frostig (2024), reveal nonlinear responses (see Section III.4.5) and differing gain estimates for the NDR and IWR read out modes. Each method produces a gain estimate within a error of the expected g=1.6 e-/DN. Furthermore, the gains vary by sensor, between the eight read out channels within each sensor, and across the pixels within those read out channels.
For simplicity, we present one measurement of the gain here and use it throughout the paper, but note the significant spread in gain by pixel, requiring careful flat-fielding. Figure 4 shows the range of slopes from a PTC for one readout channel of a WINTER sensor, taken with a set of laboratory flats on an integrating sphere. The variance is taken for a stack of ten dark-subtracted frames at each exposure time, with eight rounds of sigma clipping to reject pixels outside of three standard deviations of the mean. A line is fit to the slope of the variance versus mean signal for each pixel. This analysis predicts a mean gain of g=2.5 e-/DN but shows a spread of responses across pixels within this read out, with some pixels consistent with the designed gain of g=1.6 e-/DN. Given this uncertainty, we present all analyses throughout this study in both measured counts (DN) and converted to electrons (e-).
We measured a quantum efficiency (QE) of less than 10% for the WINTER sensors, significantly lower than the designed 80%, indicating substantial losses in photon-to-electron conversion efficiency. This measurement was validated through three independent tests: (1) laboratory characterization of the full system using a controlled light source, (2) direct laboratory measurement of the sensors’ photocurrent response, and (3) on-sky calibration with reference stars. Laboratory measurements were made using a 20-inch integrating sphere illuminated with a laser driven light source from Energetiq and a series of optical and near infrared narrow bandpass filters. The absolute optical flux was determined using a silicon photodiode with a NIST-traceable responsivity calibration, allowing the quantum efficiency to be measured between 500 and 1050 nm. The test measured a peak response of approximately 5% QE, suggesting a manufacturing issue leading to deviations from the expected efficiency (Figure 6) Furthermore, in standard QE and photon transfer curve (PTC) measurements, conversion gain and QE are inherently degenerate, necessitating an independent test to pinpoint the root cause of the discrepancy.
@@ III.4.2 Quantum efficiency
To isolate the sensor photodiode array’s intrinsic response to light from any effects introduced by the ROIC or the custom readout chain, we conducted a secondary set of QE tests. These tests employed an external source measurement unit to directly supply the sensor bias voltage, allowing precise voltage control and direct photocurrent measurement on the sensor while bypassing the readout circuitry via programmable logic on the ROIC. This approach enabled a direct measurement of photocurrent response, confirming the measurement of a peak QE of (Figure 6). We define the quantum efficiency as the ratio of the number of electrons generated per incident photon. Given an input optical power (in nW) and a measured photocurrent (in µA), the QE can be expressed as:
@@
Finally, on-sky flux observations of standard stars independently measure WINTER’s full instrument and telescope throughput. Accounting for as-measured reflectivities of the mirrors and lenses, the results confirm an average sensor QE of 10 (see Frostig et al. (2024) for a detailed calculation). This analysis includes sensor serial number SCI022, which was previously in the Starboard B position in the full instrument, which has a laboratory measured QE of 9.6% (Figure 6).
Given the results of the three independent tests, the decreases QE likely arises from a manufacturing or quality control issue during hybridization, photodiode array production, or subsequent processing. The test procedures confirm the low QE is not due to deficiencies in the ROIC or amplification/read-out chain. Since the InGaAs growth and photodiode array design made use of mature processes at Teledyne-FLIR, we conclude that this particular production lot likely encountered problems in the preparation or bonding of the InGaAs layer to the silicon multiplexer—a hypothesis that could be verified through destructive testing and microscopy of a sensor from the batch.
@@ III.4.3 Dark current
WINTER is designed for sky-background-limited photometry, with TEC-based cooling at T C to reduce dark current levels below the sky background. The smaller prototype instrument employed a single-stage TEC and demonstrated background-limited performance between -C to -C, with dark current halving for every 7∘C of cooling (Simcoe et al., 2019). However, for the manufactured WINTER sensors, a combination of TEC under-scope and insufficient process control in sensor packaging resulted in decreased cooling capacity and varied final operating temperatures across the six sensors.
Figure 7 displays the dark current in the six WINTER sensors by temperature in the final integrated camera on the telescope. The sensors reach a minimum temperature between T C and T C, unable to achieve the designed cooling of T C. The temperature discrepancy arises from two key factors. First, the TEC selection process did not include realistic modeling of the sensor packaging, resulting in the use of an undersized two-stage TEC (Laird MS2-107-10-10-12-12-11-W8). Second, the final design placed an excessive heat load on the TEC through parasitic radiative coupling and thermal conduction through wire bonds from the pins. Temperature variations between sensors are due to differences in TEC bonding, vacuum seal integrity, and the temperature of each sensor board, which impacts the TEC’s hot-side temperature. The dark current decreases approximately linearly with the logarithm of the temperature down to about C, beyond which it plateaus, possibly due to thermal infrared emission from the detector housing.
@@ III.4.4 Read noise
To evaluate the read noise in the WINTER sensors we analyze the reference rows of each sensor. These reference rows, positioned outside the active 1920 × 1080 photosensitive region, are insensitive to incident radiation. As a result, any signal variations in these pixels arise predominantly from read noise without incorporating the effects of pixel response nonlinearity or photon shot noise. Figure 8 illustrates an analysis of the signal standard deviation for each reference pixel over 384 frames in a non-destructive read ramp. The median read noise for a single readout, equivalent to the current on-sky IWR mode, is approximately 24.5 DN RMS (equivalent to 61.3 e-, assuming a gain of 2.5 e-/DN), which is close to the initial project requirement of 45 e- RMS, depending on the exact gain per pixel.
By continuously reading out the sensors with non-destructive sampling, multiple exposures of the same pixel can be averaged before a reset, reducing the random electronic noise in accordance with the theoretical scaling. Figure 8 demonstrates how this technique lowers the effective read noise as frames are binned in increasing groups from two to 64. The standard deviation of the readout signal along a best-fit line decreases, with the lowest read noise reaching approximately 3 DN (7.5 e-) when averaging in groups of 64 frames.